# Tiny is not small enough: High quality, low-resource facial animation models through hybrid knowledge distillation

### v2

@@ Abstract.
Figure 1. The animations in the first row are generated by a model with 1 billion parameters that ingests the entire audio recording. The animations in the second row are generated by our model $S_{5}+$, which has 0.8 million parameters (0.08%) and uses a context window of 512ms with a latency of 64ms (future length relative to the current frame). The audio was recorded with a headset microphone.:
@@ 1. Introduction
Low latency Latency in real-time audio-driven animation has a few causes. First, the audio input often contains some amount of future audio, usually between 100 and 300 ms (Lu et al., 2021; Mu et al., 2010; Luo et al., 2014; Hong et al., 2002; Vásquez-Correa et al., 2024). The second factor is inference time, which depends on the hardware the model is running on and how much computational resources we assign to it. While we cannot optimize end-user hardware, we can try to minimize the required computational resources. In this work, we are aiming at low latency even on low-resource devices. This is opposite to other real-time facial animation models, which often do not limit computational resources or consider the speech encoder not to be part of the model, e.g. (Tang et al., 2022; Li et al., 2023). Our goal is to reduce latency such that it is not perceivable by the human eye. In the literature, there is no consensus on what the detectable latency threshold is. According to (Younkin and Corriveau, 2008), the delay between audio and visuals needs to be below 185 ms in real-time applications to be undetectable by users. Other sources report numbers as low as 90 ms to be the acceptability threshold for users (BT et al., 1998). When both audio and animations are streamed (e.g. in an online chatting application), a total delay of 200 ms is a realistic target (Websdale et al., 2018). Informed by these numbers, we set our acceptable latency to the average of  (BT et al., 1998) and  (Younkin and Corriveau, 2008), roughly 140 ms, and are aiming to minimize the latency as much as possible.
@@
In knowledge distillation, a high-performing model (the so-called teacher) is supervising the training of a smaller model (so-called student) (Gou et al., 2021). In our case, the teacher generates animations given a large corpus of audio data. This synthetic audio-animation dataset is then used to supervise the student. In other words, the teacher generates pseudo-labels to supervise the student, similar to pseudo-labels used during self-supervision (Lee et al., 2013). In the first step of our framework, we employ a strong teacher model that contains a large pre-trained speech encoder to train a small student model. Our student models consist solely of convolutional layers and feedforward layers that operate on windows of audio input. Their architecture is thus significantly different from the teacher’s architecture, which is why this step is called heterogeneous KD (Passalis et al., 2020). In the second stage, a small student serves as a second teacher that can assist in the training of even smaller student models with smaller future context windows. Because the second teacher shares a similar architecture with the students in this homogeneous KD stage, we can use feature supervision to achieve higher performance. We call this approach hybrid knowledge distillation because it is a mixture of heterogeneous and homogeneous KD. As the homogeneous teacher has learned a task-specific representation of speech, a feature supervision loss helps lower-resource students to converge to better solutions. Our hybrid KD approach overcomes the lack of high-quality training data and allows us to reduce both latency and model size while maintaining animation quality. One example of the resulting quality is shown in Figure 1, in which we compare animations generated with the teacher model (1 billion parameters and the entire audio file as the input window) with the animations generated by a student model trained with hybrid KD (0.8 million parameters and using 64 ms of future context).
@@ 2. Introduction to knowledge distillation
Input: Unlabeled audio dataset $\mathcal{D}_{U}=\{x_{i}\}_{i=1}^{N}$, teacher model $T(\cdot)$, student model $S(\cdot)$, loss function $\mathcal{L}$, training epochs $E$
1
2for epoch $e=1$ to $E$ do
3 for each sample $x\in\mathcal{D}_{U}$ do
4      Predict pseudo-label: $\hat{y}=T(x)$;
5      Predict: $\tilde{y}=S(x)$;
@@
7      Update student $S$ model parameters using gradient descent;
8
9
10
@@ 7      Update student $S$ model parameters using gradient descent;
Input: Unlabeled audio dataset $\mathcal{D}_{U}=\{x_{i}\}_{i=1}^{N}$, teacher model $T(\cdot)$, student model $S(\cdot)$, loss function $\mathcal{L}$, hyper-parameters $\alpha_{y},\alpha_{f},$ training epochs $E$
11
12Train intermediate student $S_{0}$ according to Part 1;
13
14for epoch $e=1$ to $E$ do
15 for each sample $x\in\mathcal{D}_{U}$ do
16      Predict pseudo-label: $\hat{y}=T(x)$;
17      Predict: $\tilde{y}=S(x)$;
18      Extract intermediate features: $\hat{f}=F^{S_{0}}(x)$ and $\tilde{f}=F^{S}(x)$;
@@
20      Update student $S$ model parameters using gradient descent;
21
22
23
@@ 2.1. Knowledge distillation with pseudo-labels
As mentioned in the introduction, one important factor limiting the development of efficient facial animation models is the lack of a large dataset of paired high-quality audio-animation data. That is, relying solely on a small labeled dataset to train efficient models is far from sufficient to achieve high performance. One way to overcome the lack of labels is to create synthetic or pseudo-labels. For example, in self-supervision, a machine learning model is partially trained on labels that the model itself has predicted for unlabeled data points (Lee et al., 2013). We adopt these pseudo-labels in the setting of knowledge distillation. As detailed in the first part of Algorithm 1, we use the teacher model to generate the pseudo-labels, in our case animation frames, for a large, diverse speech dataset, de facto synthesizing a large dataset of paired high-quality audio - animation data  [^1] We were not able to find related work with similar approaches other than (Gandhi et al., 2023) who use a similar idea when distilling a large automatic speech recognition model. . This dataset is then used to train student models.

### v3

@@ Abstract.
Figure 1. The animations in the first row are generated by a model with 1 billion parameters that ingests the entire audio recording. The animations in the second row are generated by our model $S_{5}+$, which has 0.8 million parameters (0.08%) and uses a context window of 512ms with a latency of 64ms (future length relative to the current frame). The audio was recorded with a headset microphone.
Teaser
@@ 1. Introduction
Low latency Latency in real-time audio-driven animation has a few causes. First, the audio input often contains some amount of future audio, usually between 100 and 300 ms (Lu et al., 2021; Mu et al., 2010; Luo et al., 2014; Hong et al., 2002; Vásquez-Correa et al., 2024). The second factor is inference time, which depends on the hardware the model is running on and how much computational resources we assign to it. While we cannot optimize end-user hardware, we can try to minimize the required computational resources. In this work, we are aiming at low latency even on low-resource devices. This is opposite to other real-time facial animation models, which often do not limit computational resources or consider the speech encoder not to be part of the model, e.g. (Tang et al., 2022; Li et al., 2023). Our goal is to reduce latency such that it is not perceivable by the human eye. In the literature, there is no consensus on what the detectable latency threshold is. According to (Younkin and Corriveau, 2008), the delay between audio and visuals needs to be below 185 ms in real-time applications to be undetectable by users. Other sources report numbers as low as 90 ms to be the acceptability threshold for users (BT and others, 1998). When both audio and animations are streamed (e.g. in an online chatting application), a total delay of 200 ms is a realistic target (Websdale et al., 2018). Informed by these numbers, we set our acceptable latency to the average of  (BT and others, 1998) and  (Younkin and Corriveau, 2008), roughly 140 ms, and are aiming to minimize the latency as much as possible.
@@
In knowledge distillation, a high-performing model (the so-called teacher) is supervising the training of a smaller model (so-called student) (Gou et al., 2021). In our case, the teacher generates animations given a large corpus of audio data. This synthetic audio-animation dataset is then used to supervise the student. In other words, the teacher generates pseudo-labels to supervise the student, similar to pseudo-labels used during self-supervision (Lee and others, 2013). In the first step of our framework, we employ a strong teacher model that contains a large pre-trained speech encoder to train a small student model. Our student models consist solely of convolutional layers and feedforward layers that operate on windows of audio input. Their architecture is thus significantly different from the teacher’s architecture, which is why this step is called heterogeneous KD (Passalis et al., 2020). In the second stage, a small student serves as a second teacher that can assist in the training of even smaller student models with smaller future context windows. Because the second teacher shares a similar architecture with the students in this homogeneous KD stage, we can use feature supervision to achieve higher performance. We call this approach hybrid knowledge distillation because it is a mixture of heterogeneous and homogeneous KD. As the homogeneous teacher has learned a task-specific representation of speech, a feature supervision loss helps lower-resource students to converge to better solutions. Our hybrid KD approach overcomes the lack of high-quality training data and allows us to reduce both latency and model size while maintaining animation quality. One example of the resulting quality is shown in Figure 1, in which we compare animations generated with the teacher model (1 billion parameters and the entire audio file as the input window) with the animations generated by a student model trained with hybrid KD (0.8 million parameters and using 64 ms of future context).
@@ 2. Introduction to knowledge distillation
Input: Unlabeled audio dataset $\mathcal{D}_{U}=\{x_{i}\}_{i=1}^{N}$, teacher model $T(\cdot)$, student model $S(\cdot)$, loss function $\mathcal{L}$, training epochs $E$, hyper-parameters $ws,d$.
2 $\mathcal{D}_{Y}=\{\}$;
3 for each audio file $x_{i}$ do
4   Get pseudo labels $\hat{y}_{i}=\{\hat{y}_{i}^{t}\}_{t=1:t_{i}}=T(x_{i})$;
5 $\mathcal{D}_{Y}=\mathcal{D}_{Y}\ +\ \hat{y}_{i}$;
6
7
8for epoch $e=1$ to $E$ do
9 for each sample $x_{i}\in\mathcal{D}_{U}$ do
10    for frame $t\in[1,t_{i}]$ do
11         Get: $w_{t}^{d}=x_{i}[t-ws+d,t+d],\ \hat{y}=\hat{y}_{i}^{t}\in\mathcal{D}_{Y}$;
12         Predict: $\tilde{y}=S(w_{t}^{d})$;
@@
14         Update student $S$ parameters using gradient descent;
15
16
17
18
@@ 14         Update student $S$ parameters using gradient descent;
Input: Unlabeled audio dataset $\mathcal{D}_{U}=\{x_{i}\}_{i=1}^{N}$, teacher model $T(\cdot)$, student model $S(\cdot)$, loss function $\mathcal{L}$, training epochs $E$, hyper-parameters $\alpha_{y},\alpha_{f},ws,d,d^{\prime}(d^{\prime}<d).$
19
20Predict all pseudo-labels $\mathcal{D}_{Y}$ same as Part 1;
21
22Train intermediate student $S_{0}$ according to Part 1, and freeze it;
23
24for epoch $e=1$ to $E$ do
25 for each sample $x_{i}\in\mathcal{D}_{U}$ do
26    for frame $t\in[1,t_{i}]$ do
27         Get: $w_{t}^{d}=x_{i}[t-ws+d,t+d],\hat{y}=\hat{y}_{i}^{t}\in\mathcal{D}_{Y},$
28           $w_{t}^{d^{\prime}}=x_{i}[t-ws+d^{\prime},t+d^{\prime}]$;
29         Predict: $\tilde{y}=S(w_{t}^{d^{\prime}})$;
30         Extract intermediate features:
31           $\hat{f}=F^{S_{0}}(w_{t}^{d})$,  $\tilde{f}=F^{S}(w_{t}^{d^{\prime}})$;
@@
33         Update student $S$ parameters using gradient descent;
34
35
36
(In practice, we use stochastic gradient descent for both parts.)
@@ 2.1. Knowledge distillation with pseudo-labels
As mentioned in the introduction, one important factor limiting the development of efficient facial animation models is the lack of a large dataset of paired high-quality audio-animation data. That is, relying solely on a small labeled dataset to train efficient models is far from sufficient to achieve high performance. One way to overcome the lack of labels is to create synthetic or pseudo-labels. For example, in self-supervision, a machine learning model is partially trained on labels that the model itself has predicted for unlabeled data points (Lee and others, 2013). We adopt these pseudo-labels in the setting of knowledge distillation. As detailed in the first part of Algorithm 1, we use the teacher model to generate the pseudo-labels, in our case animation frames, for a large, diverse speech dataset, de facto synthesizing a large dataset of paired high-quality audio - animation data  [^1] We were not able to find related work with similar approaches other than (Gandhi et al., 2023) who use a similar idea when distilling a large automatic speech recognition model. . This dataset is then used to train student models.